OpenStudy (anonymous):

find rectangular coordinates for each point with the given polar coordinates (-1,pi/2)

myininaya (myininaya):

we want to find (x,y) we know r=-1 and theta=pi/2 we also know that x=rcos(theta) and y=rsin(tetha)

myininaya (myininaya):

so x=-1cos(pi/2)=-1(0)=0 y=-1sin(pi/2)=-1(1)=-1

myininaya (myininaya):

so (x,y)=(0,-1)